order : fentanyl 2 mg /hr . the bag is labeled 250 mg in 500 ml of solution. what is the drop factor if tubing is rated at 20 gtt/ml?
Answer #1

ANSWER

* Calculation of drop rate (gtt/min).

      Formula

          Drop rate(gtt/min) = Volume(ml) / Time(min) * Drop factor (gtt/ml).

           Calculation of volume(ml).

                Volume(ml) = Desired dose/Dose in hand * Quantity.

       Here

           Desired dose(D) = 2 mg.

           Dose in hand (H) = 250 mg.

                 Quantity(Q) = 500 ml.

   As per above formula

           Volume(ml) = 2/250 * 500

                              = 0.008 * 500

                              = 4 ml.

            Time(min) = 60 min (1 hr).

        Drop factor(gtt/min) = 20 gtt/ml.

             Drop rate(gtt/min) = Volume(ml) /Time(min) * Drop factor (gtt/ml).

                                            = 4/60 * 20

                                             = 0.07 * 20

                                             = 1.4 gtt/min.